one number is 6 more than another. the difference between their squares is 120. what is the number?
Answer by unlockmath(1688)

Hello,
Let x be one number and y represent the other.
We can write 2 equations from the information.
x=y+6
x^2-y^2=120
Now let's substitute to get:
(y+6)^2-y^2=120
Expand this out:
y^2 +12y + 36 - y^2 = 120
Simplfy:
12y + 36 = 120
Subtract 36 and divide by 12 to get:
y=7
x=13
Make sense?
RJ
www.math-unlock.com